OpenStudy (anonymous):

Todd currently has $5,000. What was the value of his money four years ago if he's earned 4 percent interest each year? A. $4,274.02 B. $4,113.51 C. $4,109.64 Will fan and medal

OpenStudy (anonymous):

I want to say C. But I think I missed a step. If you get it wrong I am so sorry. :(

OpenStudy (anonymous):

Its ok i will check

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it was wrong but its ok

OpenStudy (anonymous):

k

OpenStudy (texaschic101):

5000 = .04(5000) = 5000 - 200 = 4800 --- 1 yr ago 4800 - .04(4800) = 4800 - 192 = 4608 --- 2 yrs ago 4608 - .04(4608) = 4608 - 184.32 = 4423.68 -- 3 yr ago 4423.68 - .04(4423.68) = 4423.68 - 176.95 = 4246.73 -- 4 yrs ago I must be doing something wrong because I am not getting any of your answer choices...sorry